# Towards Privacy-aware Mental Health AI Models: Advances, Challenges, and Opportunities

### v1

@@ Abstract
Mental illness is a widespread and debilitating condition with substantial societal and personal costs. Traditional diagnostic and treatment approaches, such as self-reported questionnaires and psychotherapy sessions, often impose significant burdens on both patients and clinicians, limiting accessibility and efficiency. Recent advances in Artificial Intelligence (AI), particularly in Natural Language Processing and multimodal techniques, hold great potential for recognizing and addressing conditions such as depression, anxiety, bipolar disorder, schizophrenia, and post-traumatic stress disorder. However, privacy concerns, including the risk of sensitive data leakage from datasets and trained models, remain a critical barrier to deploying these AI systems in real-world clinical settings. These challenges are amplified in multimodal methods, where personal identifiers such as voice and facial data can be misused. This paper presents a critical and comprehensive study of the privacy challenges associated with developing and deploying AI models for mental health. We further prescribe potential solutions, including data anonymization, synthetic data generation, and privacy-preserving model training, to strengthen privacy safeguards in practical applications. Additionally, we discuss evaluation frameworks to assess the privacy-utility trade-offs in these approaches. By addressing these challenges, our work aims to advance the development of reliable, privacy-aware AI tools to support clinical decision-making and improve mental health outcomes.
@@ Introduction
Mental disorders are highly prevalent and represent a major cause of disability worldwide. The societal, economic, and personal impacts of mental health issues make swift diagnosis and treatment essential. Current diagnostic methods involve self-reported questionnaires and clinical interviews, while treatment typically consists of multiple therapy sessions with trained therapists. This approach requires therapists to dedicate substantial time to each patient, limiting their ability to treat a larger number of individuals. Combined with a shortage of trained therapists, this often leaves many patients undiagnosed. Additionally, completing self-reported questionnaires after every therapy session places a significant burden on patients.
These issues have driven the development of systems aimed at automating diagnosis and assisting therapists in treating mental disorders. Therapists often rely on various multimodal cues to diagnose mental illnesses. For example, depression exhibits distinct verbal and non-verbal characteristics, such as facial expressions [1, 2, 3], prosodic features [4, 2, 3], and semantic patterns [5]. Similarly, patients with anxiety often struggle to maintain eye contact [6, 7], particularly during conflict-laden conversations. Speech features are instrumental in detecting Post-Traumatic Stress Disorder (PTSD) [8, 9], while both speech and facial features are valuable for identifying Bipolar Disorder [10, 11]. Consequently, multimodal AI models capable of analyzing text, audio, and video data are being developed to automate diagnosis and support therapists in managing mental health conditions.
@@
Training such multimodal models requires multimodal data, typically obtained from recorded therapy sessions. However, collecting such data and training models face significant challenges due to privacy concerns. Data collection must comply with regulations like the General Data Protection Regulation (GDPR) [12] and the Health Insurance Portability and Accountability Act (HIPAA) [13], which prohibit releasing information that could disclose a person’s gender, age, or identity. Therapy session recordings inherently contain sensitive personal information, including patients’ voices and facial features, which could be misused for impersonation [14]. To ensure privacy, datasets are kept confidential, but many patients remain reluctant to record their sessions due to insufficient privacy guarantees for both the data and the models trained on it. As a result, available multimodal datasets are often small, which leads to biased models when used for training. Such small datasets also restrict the evaluation of the models’ generalizability and reliability. Furthermore, the trained model weights cannot be released, as they may inadvertently reveal private training data [15, 16, 17, 18]. Such privacy breaches could expose patients’ identities or enable impersonation, potentially worsening their mental health. These challenges significantly hinder the development and deployment of mental health AI models in real-world applications.
In recent years, there has been a growing interest in AI privacy. We examine privacy solutions that could be applied to develop privacy-aware AI models in the mental health domain. These solutions can be broadly categorized into two areas: (i) ensuring data privacy and (ii) ensuring model privacy. Data privacy involves modifying data to remove private information while retaining relevant mental health information[19, 20, 21, 22, 23, 24, 25]. An alternative approach is the creation of synthetic data for training mental health AI models [26, 27, 28, 29, 30]. Model privacy, on the other hand, focuses on privacy-preserving training methods, which enhance the robustness of models against malicious attacks [31, 32, 33, 34, 35, 36, 37]. However, implementing privacy protection methods often results in reduced model utility. Therefore, novel evaluation methods are essential to assess the privacy guarantees of these methods [38, 39, 23, 24, 40, 25, 41, 42, 43, 44, 32] and their impact on the model’s performance [38, 23, 24, 45, 25, 46, 47, 48, 49, 50, 32]. Both automatic [46, 47] and human evaluation [26, 50, 49, 23] approaches are employed to analyze the privacy-utility trade-off for each method.
@@
Figure 1: Potential solutions to address current privacy challenges and threats across modalities in mental health dataset creation, as well as in the development and evaluation of mental health AI models, to determine the privacy-utility trade-offs of the solutions.
@@
Privacy leakage from data includes personally identifiable information (PII) present in text transcripts and audio recordings of therapy sessions. Additionally, these recordings reveal the voices of patients and therapists, as well as the faces of patients. Malicious actors can also exploit the extracted audio and video features used in mental health diagnosis models to infer sensitive attributes, such as the patient’s age and gender.
@@
In the EU, personal information is protected under GDPR, which defines any information relating to an identified or identifiable natural person. An identifiable natural person is someone who can be identified, directly or indirectly, through an identifier such as a name, identification number, location data, online identifier, or factors specific to their physical, physiological, genetic, mental, economic, cultural, or social identity. Similarly, in the US, HIPAA safeguards individually identifiable health information, which includes details such as an individual’s name, address, birth date, Social Security number, and records of their past, present, or future physical or mental health conditions. Many of these types of personal information are frequently discussed in therapy sessions, such as where a person lives, their age, or any mental or physical health concerns they may have. While such information can be identified and removed in structured data formats like tables, therapy sessions often involve detailed personal narratives, which can inadvertently reveal sensitive information. As a result, textual transcripts and speech recordings of therapy sessions often contain PII that could be used to identify a patient. Even with anonymization of PII, they can still show identification vulnerabilities through the use of other public datasets [51].
@@
Speech data are classified as personal data under GDPR because they can reveal sensitive information about the speaker, including their identity, age, gender, health status, personality, racial or ethnic origin and geographical background [52]. Mental Health diagnosis often involves using audio features like Mel-frequency cepstral coefficients (MFCCs), Mel-spectrogram, and pitch extracted using tools like OpenSmile [53]. However, these features can inadvertently leak personal information, such as the patient’s age and gender [54]. Furthermore, MFCCs can be utilized for speech reconstruction [55], posing a risk of impersonation for both therapists and patients. Similarly, speech embeddings like Wav2Vec can enable voice conversion [56], further compromising the privacy of patients and therapists by exposing their unique vocal characteristics.
@@
Video recordings of therapy sessions often capture patients’ faces, as facial expressions and gaze during conversations are critical factors in mental health diagnosis. However, a patient’s face can directly reveal their identity, raising significant privacy concerns. Mental health models typically utilize facial features extracted from deep encoder models such as ResNet [57], or facial landmarks obtained through tools like OpenFace [58] for behavior, expression, and gaze analysis. These features, however, are susceptible to privacy breaches. Image reconstruction is feasible using features extracted by deep models such as AlexNet [59], and malicious actors can reconstruct faces from deep templates like FaceNet [60] through template reconstruction attacks [61]. Even facial landmarks can be exploited for facial reconstruction [62], potentially enabling identification and impersonation of patients.
@@
Trained models are often susceptible to leaking training data when subjected to attacks, such as membership inference attacks [15], from malicious actors. Song et al. [16] demonstrated that embedding models are particularly vulnerable to leaking membership information for infrequent training data inputs, which is especially concerning for small mental health datasets with a higher prevalence of rare data points. Similarly, Carlini et al. [17] highlighted the issue of neural networks memorizing unique training data, which can then be extracted from the trained models. Moreover, in text, private data can be leaked through context [32]. This is especially true for the mental health domain, where discussing life events can indirectly leak private data. Models are also prone to exposing user information contained in the data used for fine-tuning [18]. This poses a significant privacy challenge to releasing models trained or fine-tuned on mental health datasets, as they may inadvertently memorize and disclose sensitive patient information.
@@
Data anonymization involves removing PII in transcripts and audio recordings, voice anonymization in audio recordings and face anonymization in video recordings of therapy sessions to prevent identification and impersonation threats. Below, we outline approaches for anonymizing textual, audio, and visual data to ensure privacy while retaining essential information for mental health diagnosis.
@@
PII in therapy transcripts, such as names, addresses, and dates, poses identification risks. Named Entity Recognition (NER) models can detect PII and replace them with synthetically generated values that align grammatically and semantically [19, 20]. However, therapy conversations often indirectly reveal private information, making simple NER-based methods insufficient. LLMs, such as GPT-4, have shown promise in de-identifying text [21], though real-world application faces challenges like data leakage through APIs and previous Language Model (LM) based models show poor generalization across datasets [39, 20, 77]. Augmenting de-identification datasets with synthetic data [77, 20] and specialized strategies for transcribed spoken language [78] improve performance. Text rewriting [79] offers an alternative but remains untested for conversational data and risks obscuring linguistic cues critical for mental health diagnosis.
@@
Audio anonymization involves detecting and replacing PII in recorded sessions. Pipelines often use Automatic Speech Recognition (ASR) to transcribe audio, followed by NER-based PII detection and redaction. Approaches include replacing PII segments with silence [80], white noise or beeps [81]. But it makes speech unnatural. Therefore, a better approach is to use fictional content from the same category to replace PII and convert it to speech using text-to-speech or voice conversion [81]. However, this approach modifies the entire audio. Flechl et al. [22] proposed splicing matching audio fragments to generate the PII replacement and only modify the PII segment in the audio.
@@
Voice anonymization aims to protect speaker identity in data used for automatic speech and emotion recognition [23, 24]. Automatic speaker verification (ASV) systems use speaker representations like x-vectors for verification. Thus voice anonymization techniques include replacing speaker x-vectors with public x-vectors, although this reduces diversity as well as struggles with language change [82]. Orthogonal householder neural networks [82] tackle this by choosing suitable public x-vectors for maintaining diversity. However, speaker information is still present in pitch and audio bottleneck features [83] (a low-dimensional phonetic representation extracted from an intermediate layer of an ASR model). To address this, bottleneck features can be quantized [84] or perturbed with noise for differential privacy (DP) [85]. Features from pre-trained models like HuBERT [86] and OpenSmile [53] also require research for privacy-preserving extraction. Miao et al. [45] benchmarked the Multi-Speaker Anonymization (MSA), crucial for therapy recordings.
@@
Face anonymization prevents identification through video recordings. Tools like Face-Off [87], LowKey [88], Foggysight [89], and FAWKES [90] obfuscate faces in images but may fail against adaptive face recognition systems [91]. AI stylization [92] provides another alternative for face obfuscation in images while maintaining emotions of the person, crucial for mental health applications. Face anonymization in videos can be performed through applying image face obfuscation methods in every frame. However, these will be costly; therefore, specialized video face anonymization methods like FIVA [93] are more suited for mental health datasets. Other methods of image obfuscation include extracting identity representation from the image, adding noise for DP guarantees and reconstructing the image [94]. Although DP-based methods for video anonymization focus on object indistinguishability to protect human identity [95], its direct applicability for preventing facial recognition in therapy videos is unclear. These methods can also introduce demographic biases [41]. While useful attributes like emotion detection remain unaffected by obfuscation, detection of sensitive traits such as age and gender are also unaffected [25, 96], necessitating targeted anonymization methods for mental health applications.
@@
Textual synthetic data generation includes generating therapy transcripts with multi-turn dialogues. This is addressed by datasets like SoulChat [26] and SMILE [97], generated by converting single-turn psychological Q&A into multi-turn conversations via ChatGPT. CPsyCoun [47] used LLMs to generate multi-turn dialogues from counseling reports. Wu et al. [27] employed ChatGPT for zero-shot and few-shot generation of PTSD interview transcripts, improving PTSD diagnosis when combined with real datasets. SAPE [98] used genetic algorithms for creating better prompts to enhance synthetic therapy data generation. Role-playing setups, like those in Patient-$\Psi$ [28] and CACTUS [29], simulate patient-psychiatrist interactions by incorporating cognitive models and contextual details, improving realism and utility. Other synthetic text generation methods give theoretical guarantees using DP with language models [99, 100].
@@
Given the superior performance of multimodal models in mental health diagnosis, synthetic multimodal data generation is critical. Mehta et al. [49] proposed a unified framework for speech-gesture synthesis using text input, complementing textual generation methods. Style-Talker [101] integrates speech styles and chat history to generate conversational responses, supporting simulations of patient-psychiatrist dialogues in text and audio. ConvoFusion [50] adds gesture generation from text and audio, enabling text, audio, and video simulation. However, the sequential generation of modalities introduces cumulative noise and computational inefficiencies. Ng et al. [102] developed a method for generating photo-realistic avatars with gestures for dyadic conversations, addressing single-turn limitations but still relying on sequential modality generation. Chu et al. [30] created synthetic patients for medical training, producing video outputs by combining GPT-4, text-to-speech, and video generation models. While promising, these methods remain computationally intensive and lack specific applications for mental health diagnosis.
@@
Differential Privacy [103] provides theoretical privacy guarantees and is widely used to train privacy-aware models through Differentially-Private Stochastic Gradient Descent (DP-SGD) [31]. However, DP-SGD often suffers from poor performance in language modeling tasks [35]. In mental health, contextual information can inadvertently reveal private data [32]. Context-aware DP methods [32, 33] mitigate such issues by accounting for contextual leakage during training. Fine-tuning LLMs on private mental health data requires differentially private fine-tuning techniques [35, 34]. Beyond text, DP methods can be applied to conformer-based encoders for audio [36] and to models like ResNet for image and video data [37].
@@
Federated Learning (FL) is another popular privacy-preserving training method where training is distributed and locally trained model gradients are communicated to a central server [104]. This is useful for combining mental health data from different medical institutes. However, FL, on its own, provides limited privacy and is vulnerable to attacks [105, 106, 107] and leaking data through the local model weights and gradients [108]. Therefore, it is often combined with Local Differential Privacy (LDP) to improve privacy guarantees [108, 109].
@@
Autoencoders are commonly used in speech models to extract latent representations containing linguistic and paralinguistic information while obfuscating speaker identity. These models are trained to maximize downstream task performance, such as mental health prediction while minimizing speaker classification accuracy. Ravuri et al. [110] demonstrated the use of autoencoders to retain depression severity prediction performance while reducing speaker classification accuracy. Similarly, Pranjal et al. [111] used autoencoders to transform physiological, acoustic, and daily life measurements for anxiety detection on the TILES 2018 dataset while reducing identification risks.
@@
Ensuring privacy in AI models for mental health diagnosis is essential to protect patient confidentiality. However, this often comes at the cost of reduced performance in downstream diagnostic tasks. This section discusses methodologies for evaluating privacy-utility trade-offs across three key areas: data anonymization, synthetic data generation, and privacy-aware training.
@@
Privacy of PII detection and removal methods used for therapy transcripts is evaluated using standard metrics such as precision, recall, accuracy and F1-score [112, 19, 39, 77, 21, 20], to measure the ability to classify PII words. However, indirect PII leakage in therapy sessions necessitates testing against adversarial re-identification models [38], which can be enhanced using LLMs for improved privacy evaluation. It is also important to test vulnerabilities that arise from any related public dataset [51]. The privacy of PII removal methods for audio recordings is similarly measured using PII detection metrics [80, 81, 22]. Voice anonymization techniques are evaluated using the Equal Error Rate (EER) [23, 24, 82, 84, 85]or False Accept Rate (FAR) [45] of ASV systems, where higher EER and lower FAR indicate better privacy. Robustness against attack models [23, 24, 82] and unlinkability [85, 40] further assess privacy capabilities. For face anonymization in video recordings, privacy is tested by evaluating face recognition systems against anonymized faces [87, 88, 89, 90, 93], along with leakage of attributes like age and gender [25]. Robustness against facial reconstruction attack should also be tested [93]. Demographic fairness is crucial, as anonymization methods may disproportionately affect certain groups [41]. Finally, multimodal data can exacerbate privacy risks. Thus, multimodal re-identification models are essential for holistic privacy evaluation across text, audio, and video.
@@
Utility evaluation of PII removal in text measures model performance on downstream tasks using PII-removed transcripts. Sanchez et al. [112] assessed utility by calculating the proportion of information preserved, while Morris et al. [38] used metrics like masked word percentage and information loss. For therapy transcripts, the utility should focus on preserving mental health-relevant information. For PII-removed audio, utility is evaluated using metrics like substitution, hallucination, and omission percentages [81]. Such calculations should be limited to mental health-relevant segments. Additional human evaluations measuring naturalness, style consistency, and relevance should also be performed. Finally, models trained on PII-removed audio should be tested on mental health-related tasks to gauge utility. Voice anonymization utility is assessed through intelligibility (via Word Error Rate [23, 24, 82, 45]), emotion preservation (using emotion recognition performance [24]), intonation preservation (via pitch correlation [23]), and diversity (Gain of Voice Distinctiveness [23, 82]). Human evaluations of naturalness and intelligibility [23], as well as automatic measures like Predicted Mean Opinion Score (PMOS) [45], further refine utility assessment. Utility should also evaluate models trained on anonymized voices. Since therapy data is multi-speaker, multi-speaker anonymization requires utility evaluation through Diarization Error Rate (DER) [45]. For face anonymization, utility is tested by evaluating anonymized videos on downstream tasks, such as emotion detection or mental health diagnosis [25, 92].

### v2

@@ Abstract
Mental health disorders create profound personal and societal burdens, yet conventional diagnostics are resource-intensive and limit accessibility. Advances in artificial intelligence, particularly natural language processing and multimodal methods, offer promise for detecting and addressing mental disorders, but raise critical privacy risks. This paper examines these challenges and proposes solutions, including anonymization, synthetic data, and privacy-preserving training, while outlining frameworks for privacy–utility trade-offs, aiming to advance reliable, privacy-aware AI tools that support clinical decision-making and improve mental health outcomes.
@@ Introduction
Mental disorders are a leading cause of disability worldwide, with profound societal, economic, and personal consequences. Early and accurate diagnosis is critical, yet current clinical practices, relying on self-reported questionnaires and time-intensive clinical interviews, face scalability challenges due to the global shortage of trained therapists. Consequently, many individuals remain undiagnosed or untreated. Routine completion of post-session questionnaires also imposes an additional burden on patients.
These limitations have spurred interest in AI-based systems designed to automate diagnosis and assist in therapy. In clinical settings, therapists rely on a range of multimodal cues: verbal, visual, and acoustic, to assess mental health conditions. For instance, depression is associated with characteristic facial expressions [1, 2, 3], speech prosody [4, 2, 3], and semantic patterns [5]. Anxiety often manifests as difficulty maintaining eye contact [6, 7]. Speech features are instrumental in detecting Post-Traumatic Stress Disorder (PTSD) [8, 9], while both speech and facial features are valuable for identifying Bipolar Disorder [10, 11]. Accordingly, multimodal AI models [12] that integrate text, audio, and video data have shown promise in improving diagnosis and support therapists in managing mental health conditions. Recent approaches have combined acoustic features from spectrograms with visual data from convolutional neural networks [13], or incorporated facial action units, gaze, and head pose extracted using OpenFace [14] along with textual data [15]. Others utilize audio descriptors (e.g., MFCCs, eGeMAPS) together with facial landmarks, gaze, head pose and AUs to improve classification of depression and bipolar disorder [16]. Fusion methods such as adapter-based fusion [17], cross-modal attention based fusion [18] and early fusion mechanisms [19] have further improved predictions of mental disorders.
@@
Despite these advances, the development of multimodal AI for mental health remains hindered by privacy concerns. Training such models requires sensitive therapy recordings containing identifiable speech and facial data, subject to regulations including the General Data Protection Regulation (GDPR) [20] and the Health Insurance Portability and Accountability Act (HIPAA) [21]. These data cannot be publicly released, and many patients are reluctant to participate without strong privacy guarantees. As a result, existing datasets are often small and biased, limiting model performance and generalizability. Moreover, releasing model weights trained on sensitive data may inadvertently leak private information through memorization or inference attacks, posing serious ethical and legal risks [22, 23, 24, 25]. Such privacy breaches could expose patients’ identities or enable impersonation, potentially worsening their mental health. These challenges significantly hinder the development and deployment of mental health AI models in real-world applications.
To address these challenges, privacy-preserving techniques are increasingly being explored. These techniques fall into two broad categories: data privacy, which involves anonymizing while retaining clinical relevance[26, 27, 28, 29, 30, 31, 32] or synthetic data generation [33, 34, 35, 36, 37]; and model privacy, which focuses on training algorithms robust to privacy breaches, using methods such as differential privacy (DP) [38, 39, 40, 41, 42, 43, 44]. However, these protections often reduce model utility, underscoring the need for rigorous evaluation metrics for both privacy [45, 46, 30, 31, 47, 32, 48, 49, 50, 51, 39] and utility [45, 30, 31, 52, 32, 53, 54, 55, 56, 57, 39]. Both automatic [53, 54] and human [33, 57, 56, 30] evaluations are used to quantify the privacy-utility trade-off and to guide the development of effective, safe systems.
@@
Figure 1: Overview of privacy challenges, potential solutions and privacy-utility evaluations in Mental Health AI. Potential solutions to address current privacy challenges and threats across modalities in mental health dataset creation, as well as in the development and evaluation of mental health AI models, to determine the privacy-utility trade-offs of the solutions.
@@
Privacy leakage from data includes personally identifiable information (PII) present in text transcripts and audio recordings of therapy sessions. Additionally, the audio-visual therapy recordings reveal the voices of patients and therapists, as well as the faces of patients. Malicious actors can also exploit the extracted audio and video features used in mental health diagnosis models to infer sensitive attributes and PII, such as the patient’s age and gender or to reconstruct their voices and faces.
@@
In the EU, the GDPR protects PII like name, identification number, location data, online identifier, or factors specific to their physical, physiological, genetic, mental, economic, cultural, or social identity. Similarly, in the US, HIPAA protects individually identifiable health information, including details such as an individual’s name, address, birth date, Social Security Number, and records of their past, present or future physical or mental health conditions. Many of these types of personal information are frequently discussed in therapy sessions. While such information can be identified and removed in structured data formats like tables, therapy sessions often involve detailed personal narratives, which can inadvertently reveal sensitive information. As a result, textual transcripts and speech recordings of therapy sessions often contain PII that could be used to identify a patient. Even with anonymization of PII, they can still show identification vulnerabilities through the use of other public datasets [58].
@@
Speech data are classified as personal data under GDPR because they can reveal sensitive information about the speaker, including their identity, age, gender, health status, personality, racial or ethnic origin and geographical background [59]. Features commonly used in mental health models such as MFCCs, pitch, and Mel-spectrograms can unintentionally encode sensitive information such as a person’s age and gender [60]. Studies have shown that MFCCs can support speech reconstruction [61], while learned embeddings like Wav2Vec can be used for voice conversion [62], enabling impersonation of both patients and therapists.
@@
Mental health models typically utilize facial features extracted from deep encoder models such as ResNet [63] or facial landmarks, obtained through tools like OpenFace [14] for behavior, expression, and gaze analysis. However, features derived from deep learning models [64, 65] are susceptible to reconstruction attacks, enabling adversaries to recover facial identities [66, 67]. Even lower-level features like facial landmarks can be used to reconstruct approximate facial geometry [68].
@@
Trained models are often susceptible to leaking training data when subjected to attacks, such as membership inference attacks [22], from malicious actors. This is particularly concerning for models trained on mental health datasets, in which case leaks will lead to revealing private data. Mental health datasets are generally small and have a higher prevalence of rare data points. Embedding models are particularly vulnerable to leaking such infrequent training data inputs [23]. Neural networks also memorize such unique training data points, which can then be extracted from the trained models [24]. Moreover, in text, private data can be leaked through context [39]. This is especially true for the mental health domain, where discussing life events can indirectly leak private data. Models are also prone to exposing user information contained in the data used for fine-tuning [25]. This poses a significant privacy challenge to releasing models trained or fine-tuned on mental health datasets, as they may inadvertently memorize and disclose sensitive patient information.
@@
Data anonymization involves removing PII in transcripts and audio recordings, voice anonymization in audio recordings and face anonymization in video recordings of therapy sessions to prevent identification and impersonation threats. Below, we outline approaches for anonymizing textual, audio, and visual data to ensure privacy while retaining essential information for mental health diagnosis and discuss their limitations.
@@
The presence of PII in therapy transcripts, such as names, addresses, and dates, poses significant privacy risks. Common approaches employ Named Entity Recognition (NER) models to detect and replace PII with synthetic tokens [84, 26], yet these systems struggle with generalization across diverse contexts [46, 27, 85] and require domain-specific fine-tuning [28]. Existing efforts, such as dataset augmentation with synthetic data [85, 27] and methods tailored to spoken language [86], offer incremental improvements. Yet, these advances remain largely confined to Electronic Health Records (EHRs) and do not tackle the unique challenges of therapy sessions where sensitive disclosures unfold dynamically over multiple interactions. Even approaches targeting spoken data[86] rely on overly simplistic setups like passage reading, failing to capture the nuances of actual patient-therapist exchanges. Moreover, therapy conversations often contain implicit and contextualized disclosures of private information that simple NER-based systems fail to capture [87]. Recent work has explored large language models (LLMs) for de-identification [28], but their real-world utility is deeply flawed as the reliance on API-based models involves sending sensitive data to proprietary models which can be used to train these models [88] thus compromising privacy.
A particularly glaring gap is the absence of privacy frameworks designed for longitudinal therapy data. Unlike standalone medical records, therapy transcripts span multiple sessions, making cross-session privacy breaches a critical yet overlooked concern. The failure of current methods to provide theoretical DP guarantees only increases the problem, raising fundamental questions about the reliability of anonymization techniques used in mental health AI models. If we are to responsibly advance AI-driven mental health support, the field must move beyond ad hoc anonymization techniques and develop robust, theoretically grounded privacy mechanisms tailored for real-world conversational therapy data.
@@
Audio anonymization typically relies on Automatic Speech Recognition (ASR) followed by NER-based PII redaction. Simple approaches such as replacing PII with silence [89], beeps, or white noise [90] compromise naturalness and degrade usability. More advanced techniques replace PII with synthetic content from the same category, and synthesize it into speech [90]. However, this approach fundamentally alters the entire audio rather than making targeted modifications. A more refined strategy would involve selectively replacing PII segments while preserving the rest of the speech, such as splicing in matching audio fragments[29]. However, these methods are yet to be validated on longitudinal therapy data, leaving them vulnerable to cross-session linkages. Like their textual counterparts, current audio anonymization methods lack formal privacy guarantees, casting doubt on their effectiveness for real-world deployment.
@@
Voice anonymization methods, while effective for protecting speaker identity in speech and emotion recognition tasks[30, 31], remain inadequate for mental health applications. Existing techniques modify x-vectors[91], pitch, and bottleneck features[92, 93] to obscure speaker identity while preserving intelligibility and emotional cues. Some even offer theoretical DP guarantees [93]. However, these methods have not been evaluated for their ability to retain speech markers essential for mental health diagnosis, raising concerns about their applicability. An even greater limitation is their focus on Single Speaker Anonymization (SSA), whereas therapy sessions involve dynamic conversations between therapists and patients, necessitating Multi-Speaker Anonymization (MSA). While Miao et al. [52] introduced an MSA benchmark, real-world therapy scenarios involve overlapping speech and stronger attack models, which remain unaddressed. Without rigorous testing on mental health datasets and the development of DP-compliant MSA techniques, current methods fall short of ensuring privacy while preserving diagnostic speech features.
@@
Face anonymization in video recordings remains an unsolved challenge, especially for mental health applications. While tools like Face-Off[94], LowKey[95], Foggysight[96], and FAWKES[97] attempt to obfuscate faces in images, they often fail against adaptive face recognition systems[98]. AI stylization[99] offers a promising alternative by preserving emotional expressions, but its effectiveness in real-world settings remains uncertain. Frame-by-frame image anonymization is computationally expensive in video contexts. More efficient video-specific tools like FIVA [100] offer promise but introduce demographic biases [48] and may leak sensitive attributes such as age or gender [32, 101]. Even methods with DP guarantees [102, 103] remain unproven in clinical applications. Furthermore, most visual anonymization overlooks non-facial behavioral cues such as body language which may reveal gender [101]. Thus, video anonymization must go beyond facial modifications to ensure privacy without erasing clinically relevant behavioral cues.
@@
Textual synthetic data generation for therapy transcripts remains limited by the shortcomings of current methods. Approaches such as prompting LLMs[34, 104], converting single-turn psychological Q&A or counseling reports into multi-turn dialogues[33, 105, 54], and role-playing setups with integrated cognitive knowledge [35, 36] offer promising directions. Some even provide DP guarantees [106, 107]. However, these approaches typically produce short dialogues (averaging 17 turns) and fail to replicate the complexity and duration of real therapy sessions. LLMs struggle with coherence, memory, and instruction retention over extended conversations [108], limiting their clinical utility. Without a robust multi-turn generation framework that can sustain long, contextually consistent dialogues, synthetic therapy data will remain artificial and clinically inadequate.
@@
Given the superior performance of multimodal models in mental health diagnosis, synthetic multimodal data generation is critical. Current methods employ sequential pipelines to generate speech, gestures, and video from text [56, 109, 57, 110, 37], but suffer from noise accumulation, and computational inefficiencies that hinder real-world applicability. A more fundamental issue is the lack of psychiatric knowledge in these methods. Moreover, these methods remain constrained to short clips, while real therapy sessions span 30 to 60 minutes and often extend across multiple sessions. To advance synthetic data generation in mental health, future methods must support long-form, contextually coherent, multimodal generation grounded in clinical knowledge. Without such innovations, synthetic data will remain insufficient for developing clinically robust AI models.
@@
DP[111] offers strong theoretical privacy guarantees and is widely used for training privacy-aware models through Differentially-Private Stochastic Gradient Descent (DP-SGD) [38]. However, its application in language modeling remains problematic due to significant performance degradation [42]. DP techniques have been extended to conformer-based audio encoders[43] and ResNet models for image and video data[44], as well as differentially private fine-tuning methods for sensitive datasets like those in mental health[42, 41].
Yet, these efforts remain incomplete. DP methods have not been adapted for multimodal models that integrate text, audio, and video, leaving them vulnerable to cross-modal information leakage. This is concerning for multimodal mental health applications. Furthermore, therapy sessions unfold over multiple interactions, making context-aware DP essential. While recent work attempts to incorporate contextual information [39, 40], their efficacy remains untested on longitudinal mental health data. Without specialized DP techniques that account for long-range context dependencies in therapy data, existing privacy mechanisms fall short of ensuring meaningful protection in real-world mental health applications.
@@
Federated Learning (FL)[112] is widely promoted as a privacy-preserving training method, enabling decentralized model training without directly sharing patient data. Although conceptually attractive for mental health settings, FL alone offers weak privacy guarantees and remains susceptible to gradient leakage attacks [113, 114, 115], as sensitive information can still leak through model gradients and local weights[116]. Combining FL with Local Differential Privacy (LDP) improves security [116, 117], but often results in significant performance loss on realistic medical datasets, small datasets or large models [118], limiting practical adoption in mental health applications. Without substantial improvements in both privacy protection and model utility, deploying FL-based approaches in real-world mental health datasets remains premature.
@@
Autoencoders are commonly used in speech models to extract latent representations that retain mental health relevant features while obfuscating speaker identity [119, 120]. While promising in speech-based tasks, their use in multimodal therapy data remains largely unexplored.
@@
Ensuring privacy in AI models for mental health diagnosis is essential to protect patient confidentiality. However, this often comes at the cost of reduced performance in downstream diagnostic tasks. This section discusses methodologies that should be used for evaluating privacy-utility trade-offs across three key areas: data anonymization, synthetic data generation, and privacy-aware training. Without standardized, context-specific benchmarks, current approaches risk being either too lax to ensure real privacy or too restrictive to be clinically useful.
@@
Effective evaluation of privacy in mental health AI models must go beyond standard metrics such as precision, recall, or F1-score for NER-based PII removal techniques, which focus narrowly on explicit identifiers [121, 26, 46, 85, 28, 27]. Similarly, audio anonymization methods are often evaluated using PII detection metrics [89, 90, 29], but these overlook deeper privacy threats present in mental health data like indirect leaks. Advanced privacy assessment requires adversarial re-identification models, including LLM-based techniques capable of detecting indirect leaks across longitudinal therapy sessions [45]. Privacy risks must also be evaluated in relation to external datasets, where residual identifiers can re-emerge [58]. Voice anonymization techniques must be tested using Equal Error Rate (EER) [30, 31, 91, 92, 93], False Accept Rate (FAR) [52] in Automatic Speaker Verification (ASV) systems, unlinkability [93, 47] and robustness against attack models [30, 31, 91]. For video, facial anonymization must be assessed for resistance to face recognition [94, 95, 96, 97, 100] and reconstruction[100], attribute leakage (e.g., age, gender) [32], and demographic bias [48]. Body language obfuscation should also be evaluated for gender inference risks [101].
Crucially, privacy assessments remain modality-specific, ignoring the potential for cross-modal leakage, a critical issue in multimodal mental health datasets. True privacy evaluation must involve multimodal re-identification models that account for interaction between text, audio, and video.
@@
Privacy in mental health AI models is meaningless if it comes at the expense of diagnostic accuracy. Any anonymization method must preserve clinically relevant information while ensuring privacy, making utility evaluation just as critical as privacy assessment.
For text anonymization, utility should be evaluated on the preservation of clinically relevant information [121, 45] and model performance on downstream diagnostic tasks. PII removed audio utility assessments should include substitution, hallucination, and omission rates [90], particularly in speech segments crucial for mental health analysis, alongside human evaluations of naturalness and style. Models trained on anonymized audio should finally be tested for downstream task performance. Voice anonymization should be evaluated via intelligibility (Word Error Rate) [30, 31, 91, 52], emotion preservation (emotion recognition accuracy) [31], pitch correlation [30], voice diversity (Gain of Voice Distinctiveness) [30, 91]. Human evaluations (naturalness, intelligibility) [30] and automatic metrics like Predicted Mean Opinion Score (PMOS) [52] provide further insight. In case of multi-speaker anonymization situations, such as therapy sessions, the Diarization Error Rate (DER) [52] should also be evaluated to ensure that speaker separation remains intact. Finally, models using anonymized voice data should be tested on downstream mental health applications. For face anonymization, utility should be tested by evaluating the performance of anonymized videos in downstream tasks such as emotion detection and mental health diagnosis [32, 99].
However, evaluating each modality separately overlooks the bigger issue: mental health diagnosis relies on the interplay between text, audio, and video. A true test of an anonymization method’s viability is the performance of a model trained on anonymized versions of all three modalities on downstream mental health tasks. Without this, utility assessments remain incomplete, and privacy-preserving methods risk being clinically irrelevant.